OpenStudy (anonymous):

Factor. x2 – 4x + 4 A. (x – 2)2 B. (x + 4)2 C. (x + 2)2 D. (x – 4)2

OpenStudy (anonymous):

All the 2s at the end are superscripts by the way

OpenStudy (anonymous):

@iGreen @triciaal help?

OpenStudy (triciaal):

the equation is in the form ax^2 +bx + c = 0 what are the factors of a*c that add up to b

OpenStudy (triciaal):

also pay attention to the signs

OpenStudy (anonymous):

Hmmm... I haven't heard the equation before... 0.0... I just don't get how to do this in the foggiest- can you think of another way to explain it

OpenStudy (anonymous):

I'm only in 8th grade

OpenStudy (triciaal):

a, b and c are real numbers the coefficients

OpenStudy (anonymous):

Ok? So D would be the answer?

OpenStudy (triciaal):

NO I was about to show you step by step how to get the answer for any equation.

OpenStudy (anonymous):

Ok, please do then :)

OpenStudy (triciaal):

have you done this yet (a + b)^2 = a^2 +2ab + b^2?

OpenStudy (triciaal):

(a-b)^2 = a^2 -2ab + b^2

OpenStudy (triciaal):

you said you are in 8th

OpenStudy (anonymous):

Nope... ummm I can learn it though,so: the answer= x^2-4x+4

OpenStudy (anonymous):

I am in 8th grade

OpenStudy (triciaal):

back to this problem

OpenStudy (anonymous):

(x-4)^2?

OpenStudy (triciaal):

and how to get (x-2)^2

OpenStudy (anonymous):

Oh, would b=2 because √4 is 2?

OpenStudy (triciaal):

 what is a*c?